The Manchester Corporation manufactures wooden pictures frames. In order to better manage costs, the Manchester Corporation had previously developed the following standards for the manufacture of its product:

  • Each unit should have 3/4 of a pound of direct materials purchased at $12 per pound.
  • Each unit should be produced in 48 minutes at a direct labor cost of $16 per hour.

The company had the following detailed retails:

Actual production was 20,000 units using 14,600 pounds of direct materials at a total cost of $168,000 and required 11,000 direct labor hours at a total cost of $190,000.

1. What is the company's "quantity" variance related to direct materials? Use a positive number to indicate a favorable variance or a negative number to indicate an unfavorable variance.

2. What is the company's "cost" variance related to direct materials? Use a positive number to indicate a favorable variance or a negative number to indicate an unfavorable variance.

3. What is the company's total variance related to direct materials? Use a positive number to indicate a favorable variance or a negative number to indicate an unfavorable variance.
Answer #1

1) Material quantity variance = (Standard qty-actual qty)Standard price = (20000*.75-14600)*12 = 4800 F

2) Material cost variance = (Standard price-actual price)actual quantity = (12*14600-168000) = 7200 F

3) Total material variance = (20000*.75*12-168000) = 12000 F